OpenStudy (anonymous):

cos x/2=cos x. find all solutions in the interval [0,2pi).

OpenStudy (anonymous):

what angle and it's half have same cos value?

OpenStudy (anonymous):

cos (x/2) = cos (x)

OpenStudy (anonymous):

Cos[0]-1 Cos[pi/6]-sqrt[3]/2 cos[pi/4]-sqrt[2]/2 cos[pi/3]-1/2 cos[pi/2]-0

OpenStudy (anonymous):

Huh? so, is that the answer? or...

OpenStudy (anonymous):

cos(B/2) = ± sqrt([1 + cos B] / 2)

OpenStudy (anonymous):

cos[x/2]=\[\pm \sqrt{1+\cos x \over 2}\]

OpenStudy (anonymous):

how do i put that in the interval as a possible solution?

OpenStudy (anonymous):

[0,2pi)

OpenStudy (anonymous):

\[{1+\cos(x) \over 2}= \cos(x)^2\] \[{1+\cos(x) }= \ 2cos(x)^2\] (2cos(x)+1)(cos(x)-1)

OpenStudy (anonymous):

2cos(x)=-1 arcos(cos(x))=arccos(-1/2) x=2pi/3

OpenStudy (anonymous):

So it would just be 0

OpenStudy (anonymous):

the solution would be x=0,4pi/3 x=0 1/2x=0 cos(0)= 1 cos(0)=1 x=4pi/3 1/2x=2pi/3 cos(4pi/3)=-1/2 cos(2pi/3)=-1/2

OpenStudy (anonymous):

than you. this helps a lot.